A closed organ pipe can vibrate at a minimum frequency of 500 Hz. Find the length of the tube. Speed of sound in air = 340 m/s.
 

Grade:12

1 Answers

Arun
6 years ago
 V = 340 m/s Distances between two nodes or antinodes⇒ λ/4 = 25 cm ⇒ λ = 100 cm = 1 m ⇒ n = v/λ = 340 Hz.